i do not know how to solve
6e^-1.4x=21

Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)

6e^-1.4x = 21
e^-1.4x = 3.5
ln(e^-1.4x) = ln(3.5)
-1.4x = ln(3.5)
x = -ln(3.5)/1.4

Answer by stanbon(48535)

solve
6e^-1.4x=21
x=21/[6e^(-1.4)]
x=21/(6*0.24659...)
x=21/1.47958...
x=14.19319988...
---------
Cheers,
Stan H.